OpenStudy (faiqraees):

Why for some McLaurin expansions like \(\large\rm (1+x)^{-1}, ln(1+x)\ \) the value of x cannot be greater than 1?

OpenStudy (p0sitr0n):

Because they diverge

OpenStudy (mathmate):

Look up "radius of convergence", or http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeries.aspx

OpenStudy (welshfella):

they are only valid when the series converges

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle (1-x)^{-1}=\sum_{n=1}^\infty x^n}\) \(\color{black}{\displaystyle (1+x)^{-1}=\sum_{n=1}^\infty (-x)^n}\) If \(\color{black}{\displaystyle \left|x\right|\ge1}\), then the series diverges. (An elementary geometric series ratio requirement for convergence is not satisfied under the aforementioned condition.) \(\color{black}{\displaystyle \ln(1+x)=\int (1+x)^{-1}dx=\int \sum_{n=1}^\infty (-1)^nx^n~dx=x\cdot \sum_{n=1}^\infty (-1)^nnx^{n}}\) and then again you need \(\color{black}{\displaystyle \left|x\right|<1}\) for convergence.

OpenStudy (solomonzelman):

The question about ln(x+1), however is a bit harder. When |x|<1, you are having a(n increasing) exponential function in the denominator, and n in the numerator.

OpenStudy (solomonzelman):

You can actually prove that any \(\displaystyle \sum_{n=j}^{\infty}\frac{\rm increasing~power~function}{\rm exponential~growth~function}\) will always diverge, or just abide by this rule as a common sense.